OpenStudy (anonymous):

express the given logarithm in terms of common logarithms. then approximate its value to four decimal places. log2 3.2 ?

OpenStudy (anonymous):

@bmorse

OpenStudy (anonymous):

@jdoe0001

OpenStudy (whpalmer4):

Can you think of a number which is a power of 2 which when divided by a power of 10 will give you 3.2?

OpenStudy (whpalmer4):

You can write the log of a quotient as the difference of the log of the numerator and the log of the denominator: \[\log{\frac{a}{b}} = \log a-\log b\]

OpenStudy (whpalmer4):

And you can change bases of logs with the change of base property: \[\log_b x = \frac{\log_a x}{\log_a b}\]

OpenStudy (anonymous):

so what numbers go where @whpalmer4

OpenStudy (whpalmer4):

well, first, did you think of those numbers I requested?

OpenStudy (whpalmer4):

Powers of 2: 2 4 8 16 32 64 128 256 etc. Powers of 10: 10 100 1000 etc. Anything jumping out at you?

OpenStudy (anonymous):

no?? @whpalmer4

OpenStudy (whpalmer4):

Find a number on the first list (powers of 2) which can be divided by a number on the second list (powers of 10) and yield the result 3.2.

OpenStudy (whpalmer4):

The goal here is to find numbers which have "easy" logs to compute in either base 2 or base 10.

OpenStudy (bmorse):

I think you'll take the log2/log3.2 =0.59592202035

OpenStudy (bmorse):

I mean log3.2/log2, i think you switch them.

OpenStudy (bmorse):

I get 1.67807190511 or 1.6781

OpenStudy (whpalmer4):

\[log_2 3.2 = \log_2 \frac{32}{10} = \log_2 32 - \log_2 10 = \log_2 32 - \frac{\log_{10}10}{\log_{10}2} = 5 - \frac{1}{0.30103} =\]

OpenStudy (whpalmer4):

\[\log_2 32 = 5\]because \[2^5=32\]\[\log_{10} 10 = 1\] because\[10^1=10\] \[\log_{10}2 \approx 0.30103\]I just happen to know that, like I know the natural and common logs of 2,3,5,7, which allows me to find the logs of many other numbers by doing simple arithmetic. For example, \(\log_{10} 8 = 3*\log_{10}2\approx 0.90309\) because \(8=2^3\)

OpenStudy (whpalmer4):

and \[\log a^b = b\,\log a\]

OpenStudy (bmorse):

http://campuses.fortbendisd.com/campuses/documents/Teacher/2009/teacher_20090408_1245.pdf Look at #9 and #10, answers are at the bottom.